OpenStudy (anonymous):

What are the solutions to the following system of equations? y = x2 + 12x + 30 8x − y = 10

OpenStudy (dogzcatz):

Steps to first one y = x2 + 12x + 30 y=14x+30 y-14x= -14x+y=30